OpenStudy (unklerhaukus):

∫sin^2(mx) from x=-π to π

OpenStudy (jamesj):

cos 2x = cos^2 x - sin^2 x = 1 - 2 sin^2 x hence sin^2 x = 1/2 . (1 - cos 2x) and therefore \[ \sin^2 (mx) = \frac{1}{2} ( 1 - 2 \cos(2mx) ) \]

OpenStudy (jamesj):

Now integrate that

OpenStudy (unklerhaukus):

\[\int\limits_{-π}^{π}Sin^2(mx)dx\]

OpenStudy (jamesj):

still there?

OpenStudy (unklerhaukus):

\[={1 \over 2} \int\limits\limits_{−π}^{π}[\cos(2mx)+1 ]dx\]

OpenStudy (jamesj):

you dropped a minus sign

OpenStudy (unklerhaukus):

\[ {1 \over 2} \int\limits\limits_{−π}^{π}[1-\cos(2mx) ]dx \]

OpenStudy (unklerhaukus):

\[={1 \over 2 } \left[ x-{\sin(2mx) \over 2m }\right]_{π}^{-π}\]

OpenStudy (unklerhaukus):

is this right so far?, i am trying to sown that the integral is equal to π for m≠0

OpenStudy (jamesj):

If m is an integer yes. Note that in your last expression you've reversed the limits of integration.

OpenStudy (unklerhaukus):

\[={1 \over 2 } \left[ x-{sin(2mx) \over 2m }\right]_{-π}^{π}\] \[={1 \over 2 } \left[ (π-{sin(2πm) \over 2m })-(-π-{sin(-2πm) \over2m})\right]\]

OpenStudy (unklerhaukus):

\[={1 \over 2}[2π−2{\sin(2πm) \over 2m} ]\] \[=π−{\sin(2πm) \over 2m}\] and the sine function will be equal to zero at multiples of 2π (m is a positive integer) hence = \[π\delta _{0m}\] (did i miss anything or make mistakes?

OpenStudy (jamesj):

I think this is ok; writing the Dirac delta function as you have here is perhaps bit over the top and not quite accurate. I would just say the integral is equal to pi , if m is an integer pi + sin(2mpi)/m , if m is not an integer Check your problem though; I'm guessing m is an integer in which case you can just get rid of the second case.

OpenStudy (unklerhaukus):

yeah m is an interger

OpenStudy (unklerhaukus):

so the integral equals 2π ; m=n=0 { π ; m=n≠0 0 ; m≠n

OpenStudy (jamesj):

Where is this parameter n coming from? I only see n in the problem.

OpenStudy (unklerhaukus):

ops. ive been doing a few similar problems that is the result for an integral of cos(nx)cos(mx) over the same limits

OpenStudy (jamesj):

I see what you're doing now. Makes sense.

OpenStudy (unklerhaukus):

back to the question i asked: If m = 0, im getting confused

OpenStudy (jamesj):

If m = n = 0, then cos(nx)cos(mx) = 1 . 1 = 1. Hence the integral is 2pi

OpenStudy (unklerhaukus):

sin (2πm)/m ~> 0/0

OpenStudy (jamesj):

\[\lim_{x \rightarrow 0} \frac{ \sin ax}{x} = a\]

OpenStudy (jamesj):

Hence the limit you wrote down is 2pi

OpenStudy (unklerhaukus):

so, sin (2πm) = m^2 in the limit of large m interesting

OpenStudy (unklerhaukus):

*small m not large m

OpenStudy (jamesj):

No, I don't know where you're getting that from.

OpenStudy (jamesj):

The "straight" limit here is that as x --> 0 then the limit of (sin x)/x is 1.

OpenStudy (jamesj):

Now when we consider sin(ax)/x, we can write this as

OpenStudy (jamesj):

\[\lim_{x \rightarrow 0} \frac{ \sin(ax)}{x} = \lim_{x \rightarrow 0} \frac{ \sin(x)}{x/a} = \lim_{x \rightarrow 0} \ a \frac{ \sin(x)}{x} = a \ \lim_{x \rightarrow 0} \frac{ \sin(x)}{x} = a . 1 = a\]

OpenStudy (jamesj):

In other words, when ax is small sin(ax) goes like ax. Or sin(m.pi.x) goes like m.pi.x

OpenStudy (unklerhaukus):

from \[=π-{\sin(2πm) \over 2m}\] the integral would become for m=0 \[=π-{(2πm) \over 2m}, \] \[=π-π\] \[=0\]

OpenStudy (jamesj):

The direct way to see this is to put m = 0 into the original integral of sin^2(mx) = sin^2(0) = 0. The integral of 0 over -pi to pi is zero.

OpenStudy (unklerhaukus):

for m≠0\[=π-{\sin(2πm) \over 2m}\] \[=π-{0 \over 2m}\] \[=π\]

OpenStudy (jamesj):

yes

OpenStudy (unklerhaukus):

Can i wright this as \[\int\limits_{-π}^{π}\sin^2 (mx) .dx\] \[=π-{sin(2πm) \over 2m}\] \[ π;{ m≠0 } \]={\[0; {m=0}\] \[=πδ_{m0}\]

OpenStudy (unklerhaukus):

Where δ_m0 is the Kronecker delta

OpenStudy (unklerhaukus):

Thankyou so much for your help JamesJ

OpenStudy (jamesj):

sure